# Kasra Arabi, R. Teal Witter, Chinmay Hegde, Niv Cohen

### v2

@@ 3.3 Analysis
There is a threshold $\tau^{\textnormal{dist}}$ so that, for all generation noises $\mathbf{z}$, inverted noises $\mathbf{z}^{\textnormal{inv}}$, and patches $i\in[k]$,
@@
An immediate consequence of the patch distance separation assumption is that we never declare an image as watermarked if its initial noise was not generated using our watermarking scheme. In practice, such unrelated images can match with a few patches; however, it is highly unlikely for them to match with the $n^{\text{match}}$ needed for our method to declare a detection.
@@
A key property of our watermarking approach is its resistance to forgeries generated from unrelated prompts. Prior watermarking methods declare an image as watermarked as long as the pattern is embedded in the initial noise and the diffusion and inverse diffusion processes remain reasonably accurate. However, this creates vulnerabilities - an adversary could take an existing watermark and apply it to an unrelated, potentially offensive, or misleading prompt.
In contrast, our approach strengthens watermark integrity by requiring that the new prompt remains semantically close to the original. This ensures that watermarks are not erroneously detected in entirely unrelated images.
@@
Consider a suspect image $\tilde{\mathbf{x}}$ produced from our watermarking scheme with initial semantic vector $\mathbf{v}$. Let $\tilde{\mathbf{v}}$ be the (possibly quite different) semantic embedding of $\tilde{\mathbf{x}}$, and $\theta\in[-90\degree,90\degree]$ be the angle between $\mathbf{v}$ and $\tilde{\mathbf{v}}$. Set $\theta^{\text{mid}}$ as the threshold between related and unrelated semantic vectors. The probability that we identify the image as watermarked is
@@
We illustrate in the example below the sharp detection thresholds Lemma 3.2 implies; specifically how watermark detection probability varies with semantic similarity between the original and a potentially modified image. We delay the proof of Lemma 3.2 to the appendix.
@@
For instance, Figure 5(c) illustrates a separation between vectors associated with the original image and those that are unrelated, occurring at a threshold of approximately $\theta^{\text{mid}}\approx 70\degree$. When our watermarking scheme is run with $\theta^{\text{mid}}=70\degree$, $n=1024$, and $b=4$, Table 1 quantifies the probability of a watermark detection. For images with angles beyond $5\degree$ of the threshold, we almost always correctly distinguish between related and unrelated watermarked images (under assumption 3.1; see Figure 4 for the empirical curve).
Table 1: Watermark Detection Probabilities. Example detection probabilities for suspect images generated by our watermarking scheme with an angle threshold of $\theta^{\text{mid}}=70\degree$, $n=1024$ patches, and $b=4$ bits. For images with angles deviating by more than $5\degree$ from the threshold, our method distinguishes between related and unrelated watermarked images.
Semantic Angle: $\theta(\mathbf{v},\tilde{\mathbf{v}})$ Detection Probability
80 $3.06\times 10^{-6}$
75 $0.0111$
70 $0.507$
65 $0.992$
60 $1.0$
4 The Cat Attack- Harmful Image Edits
Commonly evaluated forgery attacks typically aim to forge the entire image. The forged image in that case may have no semantic connection to the image (or group of images) from which the watermark was stolen. Yet, even when such attacks are not possible, the model owner’s security might still be compromised. Here, we explore one such case, where the attacker locally modifies an existing watermarked image. Adding new content to a watermarked image may pose reputational risks to the model owner, even if the rest of the image remains unchanged. Such editing attacks are expected to be most effective against watermarks that are relatively resilient to removal attacks - a watermark that is removed after small edits will typically be removed even after edits that do not change the content.
Here, we consider an attack where a cropped object is pasted into a watermarked image. We refer to it as the Cat Attack. Our watermark method can detect such edits, and declare the image non-watermarked (or watermarked and edited) in two different ways: global Semantic Similarity Detection and local Tampering Detection (both described in Section 3). In practice, we recommend starting with Semantic Similarity Detection as a faster approach that would catch most conventional attacks. When further inspection is needed, a more local Tampering Detection can be used.
@@ 5 Empirical Analysis
In this section, we evaluate the robustness of SEAL.
@@
Prior works assume that the attacker lacks access to the model weights (which are needed for accurate DDIM noise inversion) and that the noise used during generation cannot be forged or approximated with sufficient accuracy. Going beyond previous studies, we consider here a more challenging scenario in which the attacker has full access to the model weights and can invert the generated image using the same model that produced it. The attacker’s access to the private model is taken as an upper bound for the attacker’s capability in a forgery attack [4, 17].
In our experiment, we first generate an image using watermarked noise. We then perform an inversion with the same model to recover an approximate initial noise, which is subsequently used to generate a second image. Because the attack prompt differs from the original prompt, the semantic embedding of the image $\tilde{\mathbf{v}}$ changes to ${\mathbf{v}}_{attack}$. The detection algorithm, therefore compares the estimated noise to a reference derived from ${\mathbf{v}}_{attack}$ (and not $\tilde{\mathbf{v}}$). The noise based on ${\mathbf{v}}_{attack}$ is less likely to correlate to the pattern embedded in the image, enabling the detection algorithm to declare the image as ‘not watermarked’ and evade the attack. As can be seen in Table 3, our method uniquely provides non-trivial robustness in this setting.
@@
In this experiment, we explore the resilience of our watermarking method by pasting a cat image onto a watermarked image. The cat image is randomly resized to between 30% and 60% of the watermarked image’s dimensions and placed at a random location, as exemplified in Figure 3(a) (We use here a cat to avoid dealing with not-safe-for-work material). Unlike previous watermarking techniques, which overlook semantic content and thus fail to detect such alterations, our approach identifies modifications by observing elevated $\ell_{2}$ norms in the patches where the cat is overlaid, as shown in Figure 3(b). However, since our method hinges on counting patches with an $\ell_{2}$ similarity to the reference pattern norm below a given threshold, it may occasionally miss these changes.
To increase our robustness to such attacks, we use the Tampering Detection technique detailed in Section 3. The results in Table 2 reveal that, while our method offers some robustness even without the spatial test, integrating it significantly improves detection. Other methods are vulnerable to this type of attack, as they strive not to be easily removed by small edits. Since our robustness here is in tension with our resistance to watermark removal attacks, we next analyze our method performance against these types of attacks.
Table 2: Detection AUC Under the Cat Attack. AUC of detecting edits in generated images, as described in Section 4.
@@
Table 3: Robustness to Private Model-Based Regeneration Attack. An attacker with access to the private model weights can approximate the watermarked initial noise by inverting a watermarked image using the private model. We evaluate how accurately different methods evade the false identification of unrelated images, generated with this initial noise, as watermarked.
@@
Image Transformation Attacks.
We evaluated the robustness of SEAL  under a standard suite of image transformations (see Section 12). As shown in Figure 6, SEAL achieves an average detection rate of $0.896$ under these conditions. This is comparable to some watermarking and somewhat lower than others [25]. Yet, our method is uniquely resistant to forgery. Further enhancements, such as incorporating rotation search or sliding-window search during detection, could improve its robustness against removal.
@@
A straightforward approach for our method would be to use the feature vector from the first generated image rather than the embedding of its caption. However, as illustrated in Figure 5(a), this approach fails to yield a clear separation between related and unrelated images. Consequently, we employ the blip2-flan-t5-xl [15] model for caption generation and the paraphrase-mpnet-base-v2 [19] for deriving caption embeddings, which results in a more distinct separation as shown in Figure 5. To further enhance our method’s accuracy, we fine-tuned the embedding model using 10k pairs of related captions, leading to additional improvements ( Figure 5(c) [^1] To facilitate reproducibility and encourage further research, we made our fine-tuned caption embedding model publicly available at https://huggingface.co/kasraarabi/finetuned-caption-embedding ).
Steganalysis Attack.
We evaluate the robustness of our method against a steganalysis attack [26] that attempts to approximate the watermark by subtracting non-watermarked images from watermarked ones. As shown in App. Table 4, SEAL  maintains high performance under this attack.
@@
Our method is distortion-free at the single-image level, as the noise is drawn from a pseudo-random Gaussian distribution. Consequently, all single-image quality metrics are identical to those of non-watermarked images (see also Section 14).
@@ 6 Limitation and Discussion
Although we evaluated a stronger set of forgery attacks compared to prior work, other types of forgery attacks might still potentially compromise our watermark. For example, a highly persistent attacker might attempt to gather information about the correlation between individual initial noise patches and the image semantics. While not theoretically impossible, an attacker would face several practical limitations in carrying out such an attack. Among them, are the lack of access to the private model weights, the inherent stochasticity of the watermark, and the watermark owner’s ability to deploy multiple instances of the hash function.
@@
Our method is more vulnerable to removal attacks than some existing methods. However, we believe that a sufficiently persistent attacker can remove most current watermarks. Nonetheless, increasing the robustness of watermarks holds significant societal value. It helps reduce misleading content and makes forgery attacks more difficult, which is crucial for their practical deployment.
@@ 7 Conclusion
Table 4: Robustness of Steganalysis-Based Removal. Comparison of performance metrics (AUC for SEAL, Tree-Ring, and, WIND and bit accuracy for Gaussian Shading) under various levels of averaging.
@@ 8 Additional Related Works
10 Spatial Test
@@
11 Ablation of Number of Patches and Bits.
@@
12 Transformations for the Removal Attack
We use a standard suit of transformations, including a $75^{\circ}$ rotation, $25\%$ JPEG compression, $75\%$ random cropping and scaling (C & S), Gaussian blur using an $8\times 8$ filter, Gaussian noise with $\sigma=0.1$, and color jitter with a brightness factor uniformly sampled between 0 and 6.
@@ 13 Additional Limitations and Discussion
14 Image Quality Results

### v3

@@ 3.3 Analysis
There is a threshold $\tau^{\textnormal{dist}}$ so that, for all generation noises $\mathbf{z}$, inverted noises $\mathbf{z}^{\textnormal{inv}}$, and patches indeces $i\in[k]$,
@@
An immediate consequence of the patch distance separation assumption is that we never declare an image as watermarked if its initial noise was not generated using our watermarking scheme.
@@
A key property of our watermarking approach is its resistance to forgeries generated from unrelated prompts. Prior watermarking methods declare an image as watermarked as long as the watermarking pattern is embedded in the initial noise and the diffusion and inverse diffusion processes remain reasonably accurate. However, this creates vulnerabilities - an adversary could take an existing watermark and apply it to an unrelated, potentially offensive, or misleading prompt. In contrast, our approach strengthens watermark integrity by requiring that the new prompt remains semantically close to the original. This ensures that watermarks are not erroneously detected in entirely unrelated images. We formalize this claim the the lemma below.
@@
Consider a suspect image $\tilde{\mathbf{x}}$ produced from our watermarking scheme with initial semantic vector $\mathbf{v}$. Let $\tilde{\mathbf{v}}$ be the (possibly quite different) semantic embedding of $\tilde{\mathbf{x}}$, and $\theta\in[-90\degree,90\degree]$ be the angle between $\mathbf{v}$ and $\tilde{\mathbf{v}}$. Set $\theta^{\text{mid}}$ as the threshold between semantic vectors we deem related vs. unrelated. The probability that we identify the image as watermarked is
@@
We illustrate in the example below the sharp detection thresholds Lemma 3.2 implies. Namely, we show how the watermark detection probability varies with semantic similarity between the original and a potentially modified image. We delay the proof of Lemma 3.2 to the appendix.
@@
For instance, Figure 5(c) illustrates a separation between vectors associated with the original image and those that are unrelated, occurring at a threshold of approximately $\theta^{\text{mid}}\approx 55\degree$. When our watermarking scheme is run with $\theta^{\text{mid}}=55\degree$, $n=1024$, and $b=7$ (see Figure 7 for ablation of these parameters). Figure 4 quantifies the probability of a watermark detection: We observe near-perfect separation between related and unrelated watermarked images for angles exceeding 5°5$\degree$5 ° beyond the threshold. For comparison, Section 13.2 presents the semantic angle shift resulting from the simple insertion of different objects.
@@ 4 Empirical Analysis
In this section, we empirically evaluate the robustness of SEALto different attacks.
@@
Prior works assume that the attacker lacks access to the model weights (which are needed for accurate DDIM noise inversion) and that the noise used during generation cannot be forged or approximated with sufficient accuracy [4]. Going beyond previous studies, we consider here a more challenging scenario in which the attacker has full access to the model weights and can invert the generated image using the same model that produced it. The attacker’s access to the private model is taken as an upper bound for the attacker’s capability in practical forgery attacks [4, 22, 17].
In our experiment, we first generate an image using watermarked noise. We then perform an inversion with the same model to recover an approximate initial noise, which is subsequently used to generate a second, forged, image. Because the attack prompt differs from the original prompt, the semantic embedding of the image $\tilde{\mathbf{v}}$ changes to ${\mathbf{v}}_{attack}$. The detection algorithm, therefore compares the estimated noise to a reference derived from ${\mathbf{v}}_{attack}$ (and not from $\tilde{\mathbf{v}}$). The noise pattern derived from ${\mathbf{v}}_{attack}$ during detection is less likely to correlate to the pattern embedded in the image, enabling the detection algorithm to declare the image as not watermarked and evade the attack. As can be seen in Table 2, our method uniquely provides non-trivial robustness in this setting.
We also evaluate the Latent Forgery Attack directly [17] in Section 11.
@@
A significant practical threat to the reputation of a watermark owner arises from localized modifications that shift the semantic interpretation of a watermarked image, as opposed to producing a wholly new image. To evaluate our method’s resilience against such tampering, we introduce an evaluation we term the Cat Attack.
In this experiment, a cropped object (e.g., a cat) is pasted onto a watermarked image. The cat image is randomly resized to between 30% and 60% of the watermarked image’s dimensions and placed at a random location, as exemplified in Figure 3(a). Unlike previous watermarking techniques that may overlook semantic content, our approach is designed to detect such alterations.
As shown in Figure 3(b), the pasting of the object leads to elevated $\ell_{2}$ norms in the affected patches. Quantitative results are presented in Table 1, and comparison of different object sizes can be found in the Table in Table 4. These results reveal that while our basic detection offers some robustness, integrating the local spatial test (described in Section 3) significantly improves the detection of these edits. This demonstrates a key advantage of our method. We note that in a different setting, methods such as [35, 26] offer a strong solution for tampering detection via post-hoc watermarking. Since robustness to tampering can be in tension with resistance to removal attacks, we next analyze our method’s performance against standard removal attacks.
Regeneration Based Removal Attack.
Our method is robust to regeneration-based removal attacks [36], similarly to other initial-noise-based approaches [14, 4, 34], and it significantly outperforms classical watermarking methods (see Section 13.3).
Steganalysis Removal Attack.
We evaluate the robustness of our method against a steganalysis attack [33] that attempts to approximate the watermark by subtracting non-watermarked images from watermarked ones. As shown in Table 3, SEAL  maintains high performance under this attack.
Table 1: Detection of the Cat Attack. ROC-AUC of detecting edits in generated images, as described in Section 4.
@@
Table 2: Robustness to Private Model-Based Forgery Attack. An attacker with access to the private model weights can approximate the watermarked initial noise by inverting a watermarked image using the private model. We evaluate how accurately different methods evade the false identification of unrelated images, generated with this initial noise, as watermarked.
@@
Robustness to Image Transformations.
We evaluated the robustness of SEAL  under a standard suite of image transformations (see Section 9.3). As shown in Figure 6, SEAL achieves an average detection rate of 0.8960.8960.896 under these conditions. This is comparable to some watermarking techniques and somewhat lower than others [31]. Yet, our method provides a unique resistance to forgery. Further enhancements, such as incorporating rotation search or sliding-window search during detection (see [4]), could improve its robustness against removal attempts.
@@
A straightforward approach for our method to approximate the final image semantics to embed it in the noise would be to use the visual feature vector from the proxy-generated $\mathbf{x}^{\textnormal{pre}}$ rather than the embedding of its caption. However, as illustrated in Figure 5(a), this approach fails to yield a clear separation between related and unrelated images. Consequently, we employ the captioning and caption-embedding for deriving caption embeddings, which results in a more distinct separation as shown in Figure 5(b). To further enhance our method’s accuracy, we fine-tuned the embedding model using 10k pairs of related captions, leading to additional improvements (Figure 5(c)).
@@
Our watermarking method is distortion-free at the single-image level, since the added noise is sampled from a pseudo-random Gaussian distribution, similarly to non-watermarked image generation. As a result, all single-image quality metrics remain identical to those of non‑watermarked images (see Table 7, Figure 10).
@@ 5 Limitation and Discussion
Although we evaluated a stronger set of forgery attacks compared to previous works, other types of forgery attacks might still potentially compromise our watermark. For example, a highly persistent attacker might attempt to gather information about the correlation between individual initial noise patches and the image semantics. While not theoretically impossible, an attacker would face several practical limitations in carrying out such an attack. Among them are the lack of access to the private model weights, the inherent stochasticity of the watermark, and the watermark owner’s ability to deploy multiple instances of the hash function by using multiple secret salts.
@@
Our method is more vulnerable to removal attacks than some existing methods. However, we believe that a sufficiently persistent attacker can remove most current watermarks. Nonetheless, improving watermark robustness against forgery attacks holds significant societal value - it is essential for protecting the model owner’s reputation and, consequently, for enabling practical deployment.
@@ 6 Conclusion
Table 3: Robustness of Steganalysis-Based Removal. Comparison of performance metrics (ROC-AUC) under various levels of averaging.
@@ 7 Additional Related Works
9 Implementation Details
9.1 Key Parameters
Unless otherwise stated, the results are reported with the following parameters: number of patch matching threshold $n_{match}=12$; patch-wise matching threshold $\tau=2.3$; number of projection per noise patch: $b=7$; number of noise patches $k=1024$. All parameters were chosen to optimize the overall performance.
9.2 Spatial Test
@@
9.3 Transformations for the Removal Attack
[…]
9.4 Embedding Model Fine Tuning Process
Source Prompts and Caption Pairing.
Prompts were sampled from MS-COCO and the Stable Diffusion Prompt Dataset (Gustavosta/Stable-Diffusion-Prompts). Images were generated using Stable Diffusion v2.1 (stabilityai/stable-diffusion-2-1-base), then captioned with BLIP-2 (Salesforce/blip2-flan-t5-xl). A regeneration loop used each caption as a prompt to generate a second image, which was again captioned, yielding semantically aligned $(\text{caption}_{1},\text{caption}_{2})$ pairs. Unrelated pairs were constructed by randomly mismatching captions. This procedure yielded 10,000 caption pairs.
Fine-Tuning.
A SentenceTransformer model (paraphrase-Mpnet-base-v2) was fine-tuned using MultipleNegativesRankingLoss. Training was performed for 140 epochs with a batch size of 64 and 10% warmup steps, using the AdamW optimizer with default settings.
We provide the full implementation details of the fine-tuning process to support reproducibility and enable further research. [^3] Our fine-tuned caption embedding model is publicly available at https://huggingface.co/kasraarabi/finetuned-caption-embedding.
@@ 9.4 Embedding Model Fine Tuning Process
10 Ablation of Number of Patches and Bits
@@
11 Resilience to Latent Forgery Attacks
We evaluate SEAL under the Latent Forgery Attack [17]. This attack aims to adversarially perturb non-watermarked images such that they appear watermarked by mimicking the latent representation of an originally watermarked image. This type of attacks assumes access to at least one watermarked image and attempts to shift unrelated images into the watermarked image latent region [22].
Our experiments, conducted on 100 images, demonstrated complete robustness against such attacks. Due to its semantic binding, our watermark is closely entangled with the high-level content of the original image (used by the attacker). Therefore, it is unlikely that the watermark can be transferred to images featuring unrelated content. Beyond this, the task of forging the latent patches is itself highly nontrivial. Yet, aiming to forge noise patches that better align semantically with the original watermarked image (used during the attack) might yield greater success. We leave this direction for future research.
@@ 12 Additional Limitations and Discussion
Watermarking Without a Proxy Image.
To watermark an image directly with SimHash, we may embed the noise via post hoc diffusion inpainting (see Section 4.3 of [4]), at the expense of a modest quality decrease; SSIM = 0.768. Alternatively, one may optimize an embedding of the prompt to correlate well between a givan prompt and the caption of the resulting image (similarly to Figure 5).
Reliance on DDIM inversion.
While diffusion models are not accurately mapped back to the initial noise used to generate them, our method is based on an empirical observation: patches with small enough $\ell_{2}$ differences are almost always generated from the same seed, suggesting consistent behavior under approximate inversion. The $\ell_{2}$ distance, as a metric used to determine whether a reconstructed noise patch matches the original, yields over $9\%9$ ROC-AUC.
Prompt Inversion Attack.
An attacker may choose to use prompt inversion methods such as [21]. When an attacker approximates the prompt, our semantic safeguard may be less effective. Yet, forgery attacks often aim to harm the owner’s reputation. Our method ensures that even if a forged image is produced, its semantics remain somewhat close to the original watermarked image.
Attacking the Confusion Band.
As shown in Figure 5, at certain levels of semantic similarity, our method may confuse images that are related or unrelated to the embedded caption. However, a forger’s ability to exploit this ambiguity is inherently constrained by the semantics of the original watermarked images — They can only forge images that are sufficiently similar to those whose watermark they supposedly managed to replicate. This limitation could potentially be mitigated in the future through improved embedding models.
Watermarking Capacity.
The ability to encode a very large number of distinct watermarks, or to identify one user out of many millions possible watermark owners, may raise concerns. As previous work has shown, noise-based watermarking methods can support millions of different users by using different hash salts, while still ensuring that the key patterns remain distinguishable [14, 4].
@@
13 Additional Experiments
13.1 CatAttack Performance vs. Object Scale
We varied the size of the pasted object in the CatAttack from 10% to 40% of the image area and evaluated detection performance at each scale. Table 4 reports the ROC-AUC (%) for each object scale, showing a gradual improvement from 95.4% at a 10% scale to 98.0% at a 40% scale.
Table 4: CatAttack Detection Performance vs. Object Scale. ROC-AUC (%) for different object sizes, as detected by SEAL + Spatial Test.
Scale (of image) 10 20 30 40
ROC-AUC (%) 95.4 96.2 97.7 98.0
13.2 Effect of Insertions on Semantic Embeddings and LSH Binning
[…]
Table 5: Insertion Type Analysis. The angle between the original and edited semantic embeddings, and the ratio of changed bins.
Insert Type Cat House Human
Angle () 71.2±13.871.2$\pm$ 13.871.2 ± 13.8 60.3±19.760.3$\pm$ 19.760.3 ± 19.7 68.8±17.368.8$\pm$ 17.368.8 ± 17.3
[…]
13.3 Robustness under Regeneration Attack
We benchmark SEAL against a range of detection approaches, including both generation-time and post-hoc methods such as HiDDeN [37], Stable Signature [12], TrustMark [6], and WOUAF [18]. As shown in Table 6, SEAL maintains a 98% detection accuracy under a regeneration attack [36], outperforming prior methods.
Table 6: Robustness of watermarking methods under regeneration-based removal attacks [36]. We report the watermark detection accuracy on regenerated (attacked) images (%).
Method HiDDeN StableSig TrustMark WOUAF SEAL
Acc. (%) 47 41 5 51 98
13.4 Proxy vs. Generated Image Comparison
In Figure 9, we present a visual comparison between the proxy image $x_{\text{pre}}$, which guides the watermark placement, and the final generated image $x$, illustrating SEAL’s ability to preserve semantic intent during watermarked generation.
Figure 9: Comparison between the proxy image $x_{\text{pre}}$ (left two) and the final generated image $x$ (right two). Zoom in for clarity.